OpenStudy (anonymous):

If sin theta=3/5 and theta is in quad 2 the exact form of cos (2theta) is Is this even right?

OpenStudy (anonymous):

cos(2θ) = 1 - 2sin^2θ = 1 - 2(3/5)^2 = 7/25

OpenStudy (anonymous):

You need to use the double Angle formula of cos. Since we know sin = 3/5 it makes it a little easier for us. Use \( \huge cos(2\theta ) = 1-2\sin^2\theta =\) \( \huge 1-2(\frac{9}{25}) =\) \( \huge \frac{7}{25}\)

OpenStudy (anonymous):

So I concur with you.